I am 20 years old and my sister is 8 years old. In how many years will I be twice as old as my sister
Answer by ikleyn(52787)

.
Let x = "In how many years", the value under the question.

Then

20 + x = 2*(8+x).


It is your major equation to solve.


20 + x = 16 + 2x  ====>  x = 20 - 16 = 4.


Answer.  In 4 years.


Check.   In 2 years, you will be 24 years old, while your sister will be 8 + 4 12 years old.

         And you will be twice as old as your sister then.


----------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.